Circular No. 6750

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1997 S1 (SOHO) C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports that D. Biesecker, University of Birmingham, has discovered a comet in both C3 and C2 coronagraphic data. The peak brightness was perhaps mag ∼ 7, and there may have been a small tail in some of the C2 images. Measurements by Biesecker and St. Cyr have been reduced by G. V. Williams and are published in detail on MPEC 1997-T01, together with a Kreutz-type orbital solution by the undersigned.

1997 UT  α_{2000}  δ_{2000}
Sept.29.865  12^{h}02.^{m}4  − 3^{o}57′

CATALOGUE OF COMETARY ORBITS 1997 The twelfth edition of the Catalogue of Cometary Orbits, complete

through the entries in the Sept. 1997 batch of Minor Planet Circulars, is now available. It includes a tabulation of osculating elements for the 132 numbered periodic comets (except for seven that are lost) for the standard epochs 1997 June 1 and Dec. 18, 1998 July 6 and 1999 Jan. 22. ‘Original’ and ‘future’ reciprocal semimajor axes are given for the 308 high-quality long-period orbits. There is also a more extended list of ‘X/ comets’ for which orbits cannot be computed. The 119-page Catalogue sells for $30.00 ($40.00 for airmail delivery outside North America). There is also an e-mail edition (with just the basic orbital information) for $50.00. Payment can be made by credit card. Orders can be sent to iausubs@cfa.harvard.edu or to the FAX number or postal address at the top of this Circular.

XTE J1739−302

R. M. Hjellming, National Radio Astronomy Observatory, reports: “Very Large Array observations on Sept. 26, 30 and Oct. 1 show two candidate radio sources inside and one on the edge of the error region for XTE J1739−302 (cf. IAUC 6748). Positions are: source A, α = 17^{h}38^{m}57^{s}.8, δ = −30^{o}08′43′′ (equinox 2000.0, uncertainties 0^{s}.1 and 2′′); source B, 17^{h}38^{m}52^{s}.2, −30^{o}09′01′′; source C, 17^{h}39^{m}00^{s}.6, −30^{o}11′35′′. Sources A and C have not shown significant variations about fluxes of 4.2 and 4.5 mJy at 4.86 and 8.46 GHz; however, source B may have varied between 9 and 12 mJy at 4.86 GHz from Sept. 26 to Oct. 1, and between 4 and 8 mJy at 8.46 GHz from Sept. 30 to Oct. 1. On this basis source B is the most likely candidate for the radio counterpart of XTE J1739−302. Further observations continue to determine the variability characteristics of these sources.”

1997 October 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Brian G. Marsden